# Pump-efficient Josephson parametric amplifiers with high saturation power

### v1

@@ II.3 Numerical solutions of EOM and optimization algorithm for PAE
Figure 5: Schematic of optimization algorithm steps. (1) Start from an initial optimization range $[0,R]$ and a set of initial variables $\{K^{(0)},g^{(0)}_{3},g^{(0)}_{4},\cdots\}$. (2) Calculate the cost function corresponding to the green arrows. (3) Minimize the cost function. (4)-(5) Update the optimization range $[0,R]$ and re-optimize the amplifier’s coefficients until the optimization process stops. (6) Record the optimal values of the amplifier parameters $\{K,g_{3},g_{4},\cdots\}$.
@@ III.1 Polynomial amplifiers
In this subsection, we apply our optimization algorithm to the polynomial amplifier to find a lower bound on the maximum PAE for the class of amplifiers that we consider in Fig. 2. The goal of this subsection is to set a benchmark that we will attempt to match in the next subsection with amplifiers that can be expressed as circuits. We first focus on degenerate amplifiers where $2\omega_{\text{s}}=\omega_{\text{p}}$. We fix the relative phase $\delta$ to be $1.5\pi$. This choice tends to maximize the gain in the small signal regime. We will start by establishing how the PAE depends on the order of the polynomial by optimizing a set of amplifiers with different orders to the target gain of 20 dB. We will choose the polynomial order that balances the algorithm’s run time and amplifier performance. Next, we will fix the polynomial order and vary the target gain to determine the relation between target gain and PAE. Finally, we will switch our focus to nondegenerate amplifiers and determine the dependence of PAE on target gain.
@@
Next, we sweep the target gain of the 7th-order polynomial degenerate amplifiers from 1.2 dB to 26 dB. As depicted in Fig. 7, there is a gradual decline in the PAE, dropping from nearly 100% at 1.2 dB to about 40% at 26 dB. As before, we use multiple starting points for the optimization algorithm. In Fig. 7 we plot all the resulting PAEs. At lower target gains, different starting points generally yield similar PAEs. However, in the high gain regime, we observe that the algorithm’s output begins to fluctuate (see Fig. 7), with some amplifiers tending to be trapped in local optima, as shown in Fig. 8.
@@
We used the optimal points of the degenerate amplifiers as starting points for the optimization of the nondegenerate amplifiers. The reason for this choice, as opposed to using a large number of random starting points, was that the large computational complexity of characterizing nondegenerate amplifiers made trying a large number of starting points computationally too expensive. We remind the reader that characterizing nondegenerate amplifiers is computationally much more expensive than characterizing degenerate amplifiers because the former requires the use of the direct time integration method while the latter can be much more efficiently handled by the harmonic balance method. Even within the context of the direct time integration method, nondegenerate amplifiers require longer integration time since $T_{\min}$ is large. This is due to the slight offset of the signal frequency from half the pump frequency, which increases $\gcd(\frac{2\pi}{\omega_{\text{s}}},\frac{2\pi}{\omega_{\text{p}}})$.
For the case of degenerate amplifiers, we previously observed a steady decline of the PAE with increasing target gain (see Fig. 7). On the other hand, for the case of nondegenerate amplifiers we observe that as we sweep the target gain, the PAE does not strongly depend on the target gain except for the lowest target gain point (see Fig. 9). It is possible that this behavior, i.e. lower than expected PAE at intermediate target gain, is a consequence of our choice of starting points for our optimization procedure. To try to address this possibility, we attempted to extend the optimal solution of the lowest target gain point to higher target gains. Unfortunately, we found that this was a poor choice of starting point at higher target gains as it resulted in amplifiers with large amplitude ripple, and therefore low PAE.
At high target gain, we did find nondegenerate amplifiers with reasonable performance. Indeed, for a 20 dB amplifier, we achieved a PAE of 14% (see Fig. 9). We remind the reader that our definition of PAE only accounts for the power leaving the amplifier at the signal frequency and not the idler frequency.
@@
Figure 7: PAE versus target gain of $7$-th order polynomial degenerate amplifiers, showing that PAE generally decreases as amplifier gain increases.
@@
Figure 8: Importance of initial points to optimization results. Here, we show gain versus signal amplitude curves of 20 dB $7$-th order polynomial amplifiers, where different curves represent amplifiers that were found by optimizing from six different starting points.
@@
Figure 9: Pump efficiency as a function of target gain for $7$-th order polynomial nondegenerate amplifiers.
III.2 JPA circuits
@@
In this subsection we will examine JPA circuits which are optimized to a target gain of $G_{\text{t}}=20~{}\text{dB}$, which is a common target in quantum computing applications. We will do so in both the degenerate and nondegenerate cases. We will first provide amplifiers with the highest possible PAE, where all circuit parameters as well as the damping rate $K$ are optimized. These high PAE amplifiers, however, may prove impractical to build due to high $K$ and complexity of the resulting circuit design. We will then provide JPA designs which are more likely to be buildable with modern techniques. In addition to amplifier gain, we provide bandwidth and third order intercept point (IP3) characteristics for our nondegenerate amplifier designs.
III.2.1 Highest PAE amplifiers
We find that by using JPA circuit designs, we are able to achieve a PAE almost as high as the optimized polynomial amplifiers. Specifically, we investigate amplifiers with inductive blocks composed of RF-SQUIDs shunted by three Josephson junctions in series with an applied current bias, as well as a separate flux bias through the RF-SQUID loop, as shown in Fig. 10. By tuning circuit parameters, we can achieve 20 dB gain up to a high maximum signal power. This is possible in the degenerate case, where $\omega_{\text{p}}=2\omega_{\text{s}}$, as well as in nondegenerate cases.
@@ III.2.1 Highest PAE amplifiers
In the degenerate case, we were able to optimize our amplifiers to a PAE of 34.0%. We plot the gain of this amplifier as a function of signal power in Fig. 11. In the nondegenerate case, we get a PAE of approximately 11.8%, with the gain curve shown in Fig. 13. These circuits have parameters as listed in Tab. 2. We note that our optimization is independent of the choice of capacitance $C$, and so our circuit elements and pump power can be rescaled by a different choice of $C$. In particular, inductances (e.g., $L$ and $L_{1}$) are inversely proportional to $C$, while currents (e.g., $i_{c1},i_{c2},$ and $J_{\text{ext}}$) and pump power, $P_{\text{pump}}$, are directly proportional to $C$. Here, we choose $C=0.5$ pF and provide device parameters based on this choice.
In Fig. 12 we demonstrate how increasing damping rate allows us to optimize amplifiers for higher PAE. Here, we provide the PAE of degenerate amplifiers of various damping rates, each with 25 copies of our extended RF-SQUID, optimized to highest saturation power. The resulting PAE of each amplifier increases monotonically with damping rate.
@@
Figure 11: Gain versus input signal power of the degenerate amplifier shown in Fig. 10, with optimized circuit tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2.
Figure 12: PAE of degenerate amplifiers with circuits as shown in Fig. 10. Each amplifier is optimized with a different damping rate, demonstrating increasing PAE with higher damping rate.
Figure 13: Gain versus input signal power of the nondegenerate amplifier shown in Fig. 10, with optimized circuit tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2. The signal frequency is $\omega_{\text{s}}=125/249\omega_{\text{p}}$.
parameter degenerate nondegenerate more practical degenerate more practical nondegenerate
$P_{\text{pump}}$49.5 dBm -49.5 dBm -53.3 dBm -53.0 dBm
$n$ 25 25 10 25
$K$ 3.07 $(2\pi)$ GHz 3.08 $(2\pi)$ GHz 0.8 $(2\pi)$ GHz 0.8 $(2\pi)$ GHz
@@
$i_{\text{c1}}$ 15.6 $\mu$A 16.1 $\mu$A 0.730 $\mu$A 3.36 $\mu$A
$i_{\text{c2}}$ 9.07 $\mu$A 9.11 $\mu$A 4.67 $\mu$A 1.45 $\mu$A
$L$ 79.4 pH 82.6 pH 217 pH 53.8 pH
$L_{1}$ 83.0 pH 88.6 pH 0.4 pH 0.784 pH
$\Phi_{\text{ext}}$ 5.72 $\phi_{0}$ 5.67 $\phi_{0}$ 0.472 $\phi_{0}$ 4.15 $\phi_{0}$
$J_{\text{ext}}$ 0.100 $\mu$A 0.100 $\mu$A -5.26 $\mu$A 0.394 $\mu$A
@@
Table 2: Optimized amplifier circuit parameters. Degenerate and nondegenerate columns refer to circuits as shown in Fig. 10, while more practical degenerate and more practical nondegenerate columns refer to circuits as shown in Fig. 14. We optimize with $n=25$ in each case except the more practical degenerate amplifier. In this case it proved difficult to optimize with $n=25$ and so instead we used $n=10$.
III.2.2 More practical amplifiers
While the amplifiers above have high PAEs, there remain two main challenges in physically realizing amplifiers with these designs. First is the high damping rate of $K/2\pi\approx 3$ GHz, and second is the use of $n$-many independent current sources. Here we present a simpler circuit with lower $K$ and only a single current source.
This more practical design is illustrated in Fig. 14. Parameter values that optimize this circuit for 20 dB gain can be found in Table 2. We provide parameters for both the degenerate and nondegenerate cases, with their respective gain curves plotted in Fig. 15 and Fig. 16. The PAE of the degenerate amplifier is approximately 5.91%, while that of the nondegenerate amplifier is 2.37%.
Figure 14: Amplifier circuit with a more practical design. The circuit is composed of $n$-many RF-SQUIDs, each shunted by 3 Josephson junctions with a single overall current bias.
Figure 15: Gain versus input signal power of a degenerate amplifier with a more practical circuit as shown in Fig. 14, tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2.
Figure 16: Gain versus input signal power of a nondegenerate amplifier with a more practical circuit as shown in Fig. 14, tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2. The signal frequency is $\omega_{\text{s}}=125/249\omega_{\text{p}}$.
III.2.3 Bandwidth and IP3

### v2

@@ II.3 Numerical solutions of EOM and optimization algorithm for PAE
Figure 5: Schematic of optimization algorithm steps. (a) Start from an initial optimization range $[0,R]$ and a set of initial variables $\{\omega_{0}^{(0)},K^{(0)},g^{(0)}_{3},g^{(0)}_{4},\cdots\}$. (b) Calculate the cost function corresponding to the green arrows. (c) Minimize the cost function. (d)-(e) Update the optimization range $[0,R]$ and re-optimize the amplifier’s coefficients until the optimization process stops. (f) Record the optimal values of the amplifier parameters $\{\omega_{0},K,g_{3},g_{4},\cdots\}$.
@@ III.1 Polynomial amplifiers
In this subsection, we apply our optimization algorithm to the polynomial amplifier to find a lower bound on the maximum PAE for the class of amplifiers that we consider in Fig. 2. The goal of this subsection is to set a benchmark for amplifiers that can be expressed as circuits, which we study in the next subsection. We first focus on degenerate amplifiers where $2\omega_{\text{s}}=\omega_{\text{p}}$. We fix the relative phase $\delta$ to be $1.5\pi$. This choice tends to maximize the gain in the small signal regime. We will start by establishing how the PAE depends on the order of the polynomial by optimizing a set of amplifiers with different orders to the target gain of 20 dB. We will choose the polynomial order that balances the algorithm’s run time and amplifier performance. Next, we will fix the polynomial order and vary the target gain to determine the relation between target gain and PAE. Finally, we will switch our focus to nondegenerate amplifiers and determine the dependence of PAE on target gain.
@@
Next, we sweep the target gain of the 6th-order polynomial degenerate amplifiers from 2 dB to 26 dB. As depicted in Fig. 7, there is a gradual decline in the PAE, dropping from nearly 100% at 2 dB to about 60% at 26 dB. As before, we use multiple starting points for the optimization algorithm. In Fig. 7 we plot the resulting PAEs. At lower target gains, different starting points generally yield similar PAEs. However, in the high gain regime, we observe that the algorithm’s output begins to fluctuate, with some amplifiers tending to be trapped in local optima, as shown in Fig. 8.
@@
We used the optimal points of the degenerate amplifiers as starting points for the optimization of the nondegenerate amplifiers. The reason for this choice, as opposed to using a large number of random starting points, was that the large computational complexity of characterizing nondegenerate amplifiers made trying a large number of starting points computationally too expensive. We remind the reader that characterizing nondegenerate amplifiers is computationally much more expensive than characterizing degenerate amplifiers because the former requires the use of the direct time integration method while the latter can be much more efficiently handled by the harmonic balance method. Even within the context of the direct time integration method, nondegenerate amplifiers require longer integration time since $T_{\min}$ is large. This is due to the slight offset of the signal frequency from half the pump frequency, which increases $\text{lcm}(\frac{2\pi}{\omega_{\text{s}}},\frac{2\pi}{\omega_{\text{p}}})$.
For the case of degenerate amplifiers, we previously observed a steady decline of the PAE with increasing target gain (see Fig. 7). Conversely, for nondegenerate amplifiers, the PAE increases from about 10% at 2 dB to around 15% at 20 dB where it saturates (see Fig. 9). This counterintuitive behavior can be elucidated as follows. Consider a nondegenerate amplifier in the degenerate limit of $\delta\omega=(\omega_{\text{s}}-\omega_{\mathrm{i}})/2\to 0$. In this case, the output is a single wave whose amplitude is a combination of the signal and idler components, $A=A_{\text{s}}+A_{\mathrm{i}}$. Since the power is proportional to the square of the amplitude, the signal wave power gain for the nondegenerate amplifier is four times ($\sim$ 6 dB) smaller than that of the corresponding degenerate amplifier [^2] This factor of 4 is a consequence of our definition of PAE, rather than a fundamental disadvantage of nondegenerate amplification..
For a degenerate amplifier at signal amplitude $A_{s}$ and frequency $\omega_{\text{s}}$ with a gain $G$, the PAE is expressed as: $\eta(G)=(G-1)\omega_{\mathrm{s}}^{2}A_{\mathrm{s}}^{2}$. The PAE for the corresponding nondegenerate amplifier with signal frequency $\omega_{\text{s}}+\delta\omega$ and gain $G/4$ is:
$\eta^{\prime}\left(\frac{G}{4}\right)=\left(\frac{G}{4}-1\right)(\omega_{\mathrm{s}}+\delta\omega)^{2}A_{\text{s}}^{2}\approx\frac{G-4}{4G-4}\eta(G).$ (24)
This leads to $\eta^{\prime}\to 0$ as $G/4\to 1$. Additionally, if $\eta(G)$ is a constant, $\eta^{\prime}$ increases monotonically, approaching $\eta(G)/4$. Thus, we can also get the upper bound for the nondegenerate PAE, $\eta^{\prime}\leq 25\%$, where the equality holds when $G\to\infty$ and $\eta(G)=1$. We remind the reader that our definition of PAE only accounts for the power leaving the amplifier at the signal frequency and not the idler frequency.
@@
Figure 7: PAE versus target gain of $6$-th order polynomial degenerate amplifiers, showing that PAE generally decreases as amplifier gain increases. The kink around $G_{t}=10~{}\text{dB}$ results from the penalty imposed in the cost function (23). For small target gains, this penalty is zero and does not affect the optimization process. As the target gain increases, the penalty restricts the design from achieving high PAE in order to prevent instability of the amplifier.
@@
Figure 8: Importance of initial points to optimization results. Here, we show gain versus signal amplitude curves of 20 dB $6$-th order polynomial amplifiers, where different curves represent amplifiers that were found by optimizing from six different starting points.
@@
Figure 9: Pump efficiency as a function of target gain for $6$-th order polynomial nondegenerate amplifiers. The PAE increases from about 10% at 2 dB to around 15% at 20 dB.
III.2 Coupling of low- and high-frequency modes in JPA circuits
Our amplifier circuits are generally composed of $n$-many identical repeating elements which equally divide the phase across the amplifier, as shown in Fig. 3 and described in Eq. (18) for the case of an amplifier composed of RF-SQUIDs. However, this framework for modeling the behavior of these circuits does not consider the intermediate nodes between each pair of RF-SQUIDs, and so ignores the effects of stray capacitances at these nodes which will be present in physical circuits. The introduction of these stray capacitances allows for dynamic instability of the phase at these intermediate nodes, which can excite high-frequency plasma-like modes.
In particular, this problem arises when the current-phase relation of the repeating element in the amplifier is nonmonotonic. In this case, there may be values of the total phase across the amplifier for which multiple divisions of the phase across the $n$-many elements are allowed. With the introduction of dynamic effects at the intermediate nodes between these blocks, we are able to determine how this effect disrupts amplifier performance. To model this behavior, we extend the framework for describing amplifier circuit dynamics in Subsec. II.2.2 by considering intermediate phases $\varphi_{i}$ between each of the $n$ repeating elements, with $\varphi_{1}$ representing the phase at the port and $\varphi_{n+1}$ representing the phase below the bottom element, i.e. connected to ground. At the intermediate nodes, we apply a capacitor and resistor in parallel to ground which provide dynamics and damping. This is shown in the inset in panel (a) of Fig. 10, where we display the two RF-SQUID circuit with an intermediate node. For the case of a circuit with $n$ RF-SQUIDs, we can write the EOM for the intermediate nodes as
$0=&\left(\frac{\varphi_{i}-\varphi_{i-1}}{LC}+\frac{i_{\text{c}}}{\phi_{0}C}\sin\left(\varphi_{i}-\varphi_{i-1}+\varphi_{\text{ext}}\right)\right)\\ &-\left(\frac{\varphi_{i+1}-\varphi_{i}}{LC}+\frac{i_{\text{c}}}{\phi_{0}C}\sin\left(\varphi_{i+1}-\varphi_{i}+\varphi_{\text{ext}}\right)\right)\\ &+\frac{C_{s}}{C}\ddot{\varphi}+K_{s}\dot{\varphi_{i}},\end{aligned}$ (25)
and for the two boundary nodes as
$\begin{split}\ddot{\varphi_{1}}+K\dot{\varphi_{1}}&+\frac{\varphi_{1}-\varphi_{2}}{LC}+\frac{i_{\text{c}}}{\phi_{0}C}\sin\left(\varphi_{1}-\varphi_{2}+\varphi_{\text{ext}}\right)\\ &=2K\dot{\varphi_{\text{in}}}\end{split}$ (26)
$\varphi_{n+1}$ $=0$ (27)
where $C_{s}$ is the stray capacitance to ground and $K_{s}$ is the damping rate at the intermediate nodes.
In Fig. 10, we demonstrate how the time-domain output from an amplifier composed of two RF-SQUIDs is disrupted when the current-phase relation is nonmonotonic. As shown in panel (a), multiple solutions for the phase across each RF-SQUID are allowed when a certain value of the total phase is reached, at which point the solutions for the intermediate phase trifurcate. We see that until the nonmonotonic regime of the RF-SQUIDs is reached, the amplifier behaves as predicted by ignoring intermediate nodes. In order to consider the intermediate nodes, we add a stray capacitance and damping with $C_{s}=0.001C$ and $K_{s}=0.01K$. In panel (b), we see the time-domain trace resulting from amplifying with a weak pump such that the total phase never reaches the trifurcation point. Here, the total phase as predicted without stray capacitances matches the phase when considering stray capacitances, and the intermediate phase between the RF-SQUIDs is half the total phase as expected. In panel (c), a stronger pump causes the total phase to pass into the trifurcated regime, and at this point the high-frequency mode of the amplifier is excited. Here, the trace predicted by neglecting stray capacitances drifts from the trace considering stray capacitances. Finally, in panel (d), the pump is increased further and the trifurcation is reached sooner, again causing the high-frequency mode to be excited as soon as the amplifier enters this regime. Beyond this point, the traces obtained by considering and by ignoring stray capacitances do not agree.
This behavior persists in the long-time regime for amplifiers with nonmonotonic current-phase relations and many RF-SQUIDs, as shown in Fig. 11. We simulate the behavior with an inductive block composed of 10 RF-SQUIDs with dynamics at each intermediate node as specified by Eq. (25). We do this for RF-SQUIDs with both monotonic and nonmonotonic current-phase relations, and we find that in the monotonic case the solution converges to the solution with no stray capacitances. Conversely, in the nonmonotonic case, high-frequency modes are always excited and the solution never converges to the no stray capacitance case. We provide additional traces in short- and long-time regimes for varying degrees of monotonicity in Appendix D.
Evidence presented in this section indicates that we should focus on amplifiers with monotonic current-phase relations. We use this as a design rule for amplifier circuits in the remainder of the paper. To enforce this design rule, we follow the prescription of Appendix C.
Figure 10: Coupling of low- and high-frequency modes induced by nonmonotonic current-phase relation of constituent RF-SQUIDs. (a) Intermediate phase (i.e. the phase difference across the bottom RF-SQUID, see inset) as a function of the phase difference across both RF-SQUIDs. At the points where the current-phase relation of the RF-SQUID becomes nonmonotonic (indicated by the thin red lines), the solution of the intermediate phase equation trifurcates. Inset: circuit diagram of the amplifier with two RF-SQUIDs. Stray capacitance and damping are indicated in orange. (b-d) Real time dynamics of the amplifier with two RF-SQUIDs depicted in the inset of panel (a) for different pump strength. Panel (b) depicts the weak pump case, where the total phase narrowly avoids the nonmonotonic regime and hence the high-frequency mode of the intermediate phase is not excited. Panel (c) depicts a slightly stronger pump, where the total phase narrowly hits the nonmonotonic regime, which results in the ringing up of the high-frequency mode of the intermediate phase. As the high-frequency mode rings up the dynamic solution with the stray capacitor starts to diverges from the dynamic solution without the stray capacitor. Panel (d) depicts an even stronger pump, where the total phase hits the nonmonotonic regime earlier.
Figure 11: Long-time behavior of amplifiers with inductive blocks with differing current-phase relations. A.U. indicates arbitrary units. (a) Dynamics for an amplifier composed of 10 RF-SQUIDs with monotonic current-phase relation (see inset). The time-domain trace indicates that the phase below the first RF-SQUID and above the following 9 is nearly exactly 9/10 the total phase. Or rather, the phase is divided equally among the RF-SQUIDs and the amplifier is well modeled without stray capacitances. (b) Dynamics for an amplifier composed of 10 RF-SQUIDs with nonmonotonic current-phase relation (see inset). The time-domain trace indicates that the phase below the first RF-SQUID and above the following 9 is unstable, since the nonmonotonicity allows for several ways to divide the phase among the RF-SQUIDs. The amplifier is not well modeled when neglecting stray capacitances since the traces do not agree.
III.3 Performance of JPA circuits
@@
In this subsection we will examine JPA circuits which are optimized to a target gain of $G_{\text{t}}=20~{}\text{dB}$, which is a common target in quantum computing applications. We will do so in both the degenerate and nondegenerate cases. In both cases, we optimize with a pump frequency $\omega_{\text{p}}=12~{}(2\pi)$ GHz, with $\omega_{\text{s}}=0.5\,\omega_{\text{s}}=6~{}(2\pi)$ GHz in the degenerate case and $125/249\,\omega_{\text{p}}$ in the nondegenerate case. In optimizing our amplifiers, we treat pump power as a hyperparameter which is fixed for each optimization, and we adjust it to achieve the highest possible PAE. We will first investigate amplifiers whose inductive blocks are composed of RF-SQUIDs, where all circuit parameters as well as the damping rate $K$ are optimized. Then, we investigate amplifier designs with higher PAE by shunting the RF-SQUIDs with three Josephson junctions with an applied current bias. This extended RF-SQUID circuit offers higher PAE at the expense of additional complexity, which may prove impractical to build. In addition to amplifier gain, we provide bandwidth and third order intercept point (IP3) characteristics for our nondegenerate amplifier designs.
III.3.1 RF-SQUID amplifiers
We find that by tuning the parameters of an amplifier composed of a chain of identical RF-SQUIDs as in Fig. 3, we can achieve a PAE which is a substantial fraction of the PAE of polynomial amplifiers. In Table 2, we list the circuit parameters for our optimized amplifiers. In particular, we tune the Josephson critical current $i_{c}$, the inductance $L$, the applied flux bias $\Phi_{\text{ext}}$, and the damping rate $K$. In the degenerate case, we also tune the relative phase offset $\delta$ between the signal and pump.
@@ III.3.1 RF-SQUID amplifiers
In the degenerate case, we were able to optimize our amplifiers to a PAE of 37.9%. We plot the gain of this amplifier as a function of signal power in Fig. 12. In the nondegenerate case, we get a PAE of approximately 8.9%, with the gain curve shown in Fig. 13. These circuits have parameters as listed in Tab. 2. We note that our optimization is independent of the choice of capacitance $C$, and so our circuit elements, pump power, and characteristic impedance $Z$ can be rescaled by a different choice of $C$. In particular, inductances (e.g., $L$) are inversely proportional to $C$, while currents (e.g., $i_{c}$) and pump power, $P_{\text{pump}}$, are directly proportional to $C$. Here, we choose $C=0.5$ pF and provide device parameters based on this choice. Our amplifiers are optimized with a low relative pump power (such that $(P_{\text{pump}})/(n^{2}CK\phi_{0}^{2}\omega_{\text{p}}^{2})$ is small), as we find this improves PAE. Higher output saturation power can be achieved without changing PAE by rescaling parameters as described above.
Figure 12: Gain versus input signal power of the degenerate RF-SQUID amplifier shown in Fig. 3, with optimized circuit tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2.
Figure 13: Gain versus input signal power of the nondegenerate RF-SQUID amplifier shown in Fig. 3, tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2. The signal frequency is $\omega_{\text{s}}=125/249\,\omega_{\text{p}}$.
III.3.2 Extended RF-SQUID amplifiers
Here, we investigate amplifiers with inductive blocks composed of RF-SQUIDs shunted by three Josephson junctions in series with an applied current bias, as well as a separate flux bias through the RF-SQUID loop, as shown in Fig. 14. With these extended RF-SQUID designs, we can achieve a PAE which is higher than that of the RF-SQUID amplifiers above by more carefully tuning higher order terms in the current-phase relation of the inductive block of the amplifier.
Parameter values that optimize this circuit for 20 dB gain can be found in Tab. 2. We provide parameters for both the degenerate and nondegenerate cases, with their respective gain curves plotted in Fig. 15 and Fig. 16. The PAE of the degenerate amplifier is approximately 42.6%, while that of the nondegenerate amplifier is 8.9%.
@@ III.3.2 Extended RF-SQUID amplifiers
Figure 15: Gain versus input signal power of the degenerate extended RF-SQUID amplifier shown in Fig. 14, with optimized circuit tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2.
Figure 16: Gain versus input signal power of the nondegenerate extended RF-SQUID amplifier shown in Fig. 14, with optimized circuit tuned to 20 dB gain. Parameters for this amplifier are listed in Tab. 2. The signal frequency is $\omega_{\text{s}}=125/249\,\omega_{\text{p}}$.
parameter degenerate RF-SQUID nondegenerate RF-SQUID degenerate extended RF-SQUID nondegenerate extended RF-SQUID
$P_{\text{pump}}$72.3 dBm -72.4 dBm -72.2 dBm -72.1 dBm
$n$ 10 10 10 10
$K$ 2.42 $(2\pi)$ GHz 2.36 $(2\pi)$ GHz 2.50 $(2\pi)$ GHz 2.57 $(2\pi)$ GHz
@@
$Z$ 131 $\Omega$ 135 $\Omega$ 127 $\Omega$ 124 $\Omega$
$i_{\text{c}}$ 18.8 $\mu$A 18.1 $\mu$A
$i_{\text{c1}}$ 21.4 $\mu$A 20.3 $\mu$A
$i_{\text{c2}}$0.921 $\mu$A* -3.32 $\mu$A*
$L$ 17.5 pH 18.1 pH 15.3 pH 15.0 pH
$L_{1}$ 399 pH 66.5 pH
$\Phi_{\text{ext}}$ 3.09 $\phi_{0}$ 3.09 $\phi_{0}$ 3.03 $\phi_{0}$ 3.09 $\phi_{0}$
$J_{\text{ext}}$ 4.65 $\mu$A 0.133 $\mu$A
[…]
$PAE$ 37.9% 8.9% 42.6% 14.2%
@@
Table 2: Optimized amplifier circuit parameters. Degenerate RF-SQUID and nondegenerate RF-SQUID columns refer to circuits as shown in Fig. 3 and described in Subsec. III.3.1, while degenerate extended RF-SQUID and nondegenerate extended RF-SQUID columns refer to circuits as shown in Fig. 14 and described in Subsec. III.3.2.
*Junctions with negative Josephson critical currents should be replaced with DC-SQUIDs with appropriate flux bias which produce this equivalent negative $i_{\text{c}}$.
III.3.3 Bandwidth and IP3